# Authors & Affiliations

### v1

@@ Abstract
With the growing integration of Artificial Intelligence (AI) in educational contexts, university students preparing for English language tests increasingly alternate between traditional search engines, such as Google, and large language models (LLMs) to support their test-related problem-solving. This study examines students’ perceptions of these tools, focusing on usability, efficiency, and their integration into English language test preparation workflows. Using a mixed-methods approach, we surveyed 140 university students from diverse academic disciplines and conducted in-depth interviews with 20 participants. Quantitative analyses, including ANOVA and chi-square tests, were employed to evaluate differences in perceived efficiency, satisfaction, and overall tool preference. The qualitative findings indicate that students frequently switch between GPT and Google depending on task demands, relying on Google for credible, multi-source information and rule verification, while using GPT for summarization, explanation, paraphrasing, and drafting responses for English test tasks. As neither tool alone was found to adequately support all aspects of English language test problem solving, participants expressed a strong preference for a hybrid solution. In response, we propose a prototype in the form of a chatbot embedded within a search interface, combining GPT’s conversational strengths with Google’s reliability to improve English language test preparation and reduce cognitive load.
Keywords: Large Language Models in Education; English Language Test Preparation; Hybrid Search and Learning Environments; Human–Computer Interaction; AI-Supported Learning

### v2

# Mark Protik Mondol1, Jannatun Noor Mukta1[0000-0001-9669-151X]

@@ Abstract
As Artificial Intelligence (AI) becomes increasingly integrated into education, university students preparing for English language tests are frequently shifting between traditional search engines like Google and large language models (LLMs) to assist with problem-solving. This study explores students’ perceptions of these tools, particularly in terms of usability, efficiency, and how they fit into English test preparation practices.Using a mixed-methods design, we collected survey data from 140 university students across various academic fields and conducted in-depth interviews with 20 participants. Quantitative analyses, including ANOVA and chi-square tests, were applied to assess differences in perceived efficiency, satisfaction, and overall tool preference. The qualitative results reveal that students strategically alternate between GPT and Google based on task requirements. Google is primarily used for accessing reliable, multi-source information and verifying rules, whereas GPT is favored for summarizing content, providing explanations, paraphrasing, and drafting responses for English test tasks.Since neither tool independently satisfies all aspects of English language test preparation, students expressed a clear preference for an integrated approach. In response, this study proposes a prototype chatbot embedded within a search interface, combining GPT’s interactive capabilities with Google’s credibility to enhance test preparation and reduce cognitive load.
Keywords: Large Language Models in Education; English Language Test Preparation; Hybrid Search and Learning Environments; Human Computer Interaction; AI-Supported Learning;Search Engine
@@ 9 Ethical Considerations
Acknowledgments
We sincerely thank everyone who participated in the survey and interviews for their time and experience sharing. Also, thanks to OpenAI for providing the free version of ChatGPT, which was used to improve grammatical accuracy and language proficiency.
Funding
This research was supported by the Institution for Advanced Research Publication Grant of United International University (Reference No.: IAR-2026-Pub-000).
Competing Interests
The authors declare that they have no competing interests.